### v1

# Decoding Translation-Related Functional Sequences in 5 UTRs Using Interpretable Deep Learning Models

Abstract
Understanding how 5’ untranslated regions (5’UTRs) regulate mRNA translation is critical for controlling protein expression and designing effective therapeutic mRNAs. While recent deep learning models have shown promise in predicting translational efficiency from 5’UTR sequences, most are constrained by fixed input lengths and limited interpretability. We introduce UTR-STCNet, a Transformer-based architecture for flexible and biologically grounded modeling of variable-length 5’UTRs. UTR-STCNet integrates a Saliency-Aware Token Clustering (SATC) module that iteratively aggregates nucleotide tokens into multi-scale, semantically meaningful units based on saliency scores. A Saliency-Guided Transformer (SGT) block then captures both local and distal regulatory dependencies using a lightweight attention mechanism. This combined architecture achieves efficient and interpretable modeling without input truncation or increased computational cost. Evaluated across three benchmark datasets, UTR-STCNet consistently outperforms state-of-the-art baselines in predicting mean ribosome load (MRL), a key proxy for translational efficiency. Moreover, the model recovers known functional elements such as upstream AUGs and Kozak motifs, highlighting its potential for mechanistic insight into translation regulation.
Introduction
Messenger RNA (mRNA) therapeutics have emerged as a powerful platform for precise control of protein expression, with broad applications in vaccine development, cancer immunotherapy, and treatment of genetic disorders (Qin et al. 2022). A critical determinant of therapeutic efficacy is translation efficiency, which governs protein yield from synthetic mRNA constructs. The 5’ untranslated region (5’UTR) plays a central role in regulating translation initiation(Araujo et al. 2012), with sequence features such as upstream AUGs (uAUGs) (Wang and Rothnagel 2004) and Kozak (Li et al. 2017) motifs modulating ribosome scanning and start-site selection. Understanding the regulatory code embedded in 5’UTRs is thus essential for rational mRNA design (Cao et al. 2021).
Recent advances in high-throughput experimental techniques, including ribosome profiling (Brar and Weissman 2015) and systematic mutagenesis (Kang et al. 2004), have produced large-scale datasets linking 5’UTR sequences to translational outcomes. These resources have enabled the application of deep learning to model translation regulation (Zhang et al. 2017), with mean ribosome load (MRL) frequently used as a key quantitative proxy for translation efficiency. Despite encouraging progress, current models face two critical limitations (Sample et al. 2019).
First, most architectures are trained on fixed-length inputs and lack the capacity to generalize to native 5’UTRs of varying lengths, requiring truncation that may obscure important regulatory elements. Second, existing models offer limited interpretability. Although some incorporate attention mechanisms or structural priors, they typically fail to attribute functional outcomes to specific sequence motifs such as uAUGs or Kozak elements (Kozak 1987), restricting their utility in mechanistic inference and variant prioritization.
In addition, current models often struggle to scale in multi-task settings or to operate efficiently in resource-constrained environments, limiting their applicability in complex biological systems and real-world deployment scenarios.
To address these challenges, we present UTR-STCNet, a Transformer-based deep learning framework for predicting translational efficiency from 5’UTR sequences. UTR-STCNet introduces two key innovations: a Saliency-Aware Token Clustering (SATC) module that groups and filters sequence tokens based on regulatory relevance, and a Saliency-Guided Transformer (SGT) block that models motif interactions across both local and long-range contexts using sparse attention. Our design enables accurate modeling of variable-length 5’UTRs with strong biological interpretability and computational efficiency. Our contributions are summarized as follows:
  • We propose UTR-STCNet, a deep learning framework for predicting mean ribosome load from 5’UTR sequences, with strong generalization across datasets and biological contexts.
  • The model supports variable-length inputs, scales to multi-task settings, and maintains a lightweight architec- ture suitable for large-scale or resource-constrained de- ployment.
  • UTR-STCNet offers intrinsic interpretability by explic- itly identifying regulatory motifs such as uAUGs and Kozak sequences.
  • UTR-STCNet consistently outperforms existing baselines across tasks involving multiple species and cell lines, achieving state-of-the-art performance in translation prediction.
Related Work
Self-Supervised Representation Learning for RNA Sequences
A growing body of work has focused on learning contextual representations of RNA sequences through self-supervised pretraining, with the goal of capturing transcript-level regulatory information. Early methods relied on handcrafted k-mer features or shallow encoders, which lacked the capacity to model long-range dependencies and complex structural signals(Alipanahi et al. 2015). To address these limitations, RNA-specific language models such as RNA-BERT (Akiyama and Sakakibara 2022) and RNA-FM (Chen et al. 2022) were introduced, trained on large-scale transcriptomic datasets to produce contextualized RNA embeddings. These models significantly improved the modeling of splicing signals, secondary structure contexts, and untranslated region (UTR) motifs. In parallel, general-purpose contextual learning frameworks such as data2vec (Baevski et al. 2022) have been adapted to RNA domains, offering a unified modality-agnostic representation space. Building on this direction, mRNA2vec (Zhang et al. 2025) proposes a joint encoding strategy for 5’UTRs and coding regions by integrating both contextual and structural cues, enabling holistic modeling of mRNA molecules. Although genomic models like DNABERT have shown strong performance in DNA sequence tasks, their architecture and training focus limit their applicability to RNA-specific regulatory interpretation. Recent multimodal extensions incorporating structural or epigenomic data further expand the landscape of transferable RNA representations.
Supervised Deep Learning for 5’UTR-Based Translation Modeling
In contrast to general-purpose representation learning, another line of work has focused on directly modeling translation efficiency from 5’UTR sequences using supervised deep learning. The 5’UTR plays a central role in post-transcriptional regulation, modulating translation initiation, ribosome recruitment, and start-site selection (Leppek, Das, and Barna 2018). Optimus (Sample et al. 2019) leverages global-pooling convolutional networks (CNNs) trained on mutagenesis data to predict mean ribosome load (MRL) from fixed-length sequences, capturing the functional impact of sequence variants. FramePool (Karollus, Avsec, and Gagneur 2021) extends this approach with a length-agnostic CNN architecture that enables frame-wise ribosome occupancy prediction, offering fine-resolution insights into initiation dynamics. To improve generalization across species and experimental conditions, MTrans (Zheng et al. 2023) applies a multi-task framework with shared CNN encoders and task-specific GRUs. UTR-LM (Chu et al. 2024) , in turn, pretrains a Transformer-based model on large-scale collections of 5’UTR sequences, learning transferable contextual embeddings for downstream translation-related tasks. Together, these models provide a strong foundation for 5’UTR-based translation modeling, demonstrating reliable performance across a range of predictive tasks. However, existing methods still face challenges in capturing interpretable, motif-aware dependencies and in modeling variable-length inputs without sacrificing biological relevance. These limitations motivate the development of specialized architectures that are both structurally adaptive and biologically transparent.
Method
We present the architecture of UTR-STCNet in Figure.1. The model incorporates a basic Transformer block (Vaswani et al. 2017), a Saliency Awared Token Clustering (SATC) Module, a Saliency Guided Transformer (SGT) Block, and a multi-task prediction heads. SATC module hierarchically merges relevant tokens into saliency-aware representatives, enhancing structural compactness and semantic coherence. These refined tokens are subsequently processed by the SGT, which compresses the representations of 5’UTRs into more compact, saliency-weighted embeddings, thereby improving computational efficiency and reducing model complexity while preserving the ability to extract critical regulatory features and capture long-range dependencies. We adopt a multi-task learning framework where the feature extraction backbone is shared across all tasks. To capture task-specific characteristics, we design independent prediction heads for each task. Each task-specific head consists of a lightweight feedforward network that maps the shared representation to the task-specific output space. The task-specific heads then predict mean ribosome load based on the aggregated and interpretable representations.
Problem Formulation
We study the problem of predicting translation-associated activity from genomic sequences. Given a DNA sequence of length $L$, denoted as $\mathbf{s}=(s_{1},s_{2},\ldots,s_{L})$ where each $s_{i}\in\{\mathrm{A},\mathrm{C},\mathrm{G},\mathrm{T}\}$, our goal is to learn a mapping function $f_{\theta}:\mathcal{S}\rightarrow\mathbb{R}_{\geq 0}$ where $\mathcal{S}$ denotes the space of all possible DNA sequences of length $L$, that estimates the MRL. The prediction is denoted as $\hat{y}_{\mathrm{MRL}}=f_{\theta}(\mathbf{s})$.
To ensure consistent input representation, each sequence $\mathbf{s}$ is embedded into a matrix $\mathbf{X}\in\mathbb{R}^{L\times d}$, where $d$ is the dimensionality of the token embeddings based on nucleotides. The labeled training dataset is denoted as $\{(\mathbf{s}^{(i)},y_{\mathrm{MRL}}^{(i)})\}_{i=1}^{N}$, where $N$ is the number of training samples.
Figure 1: Overview of UTR-STCNet. The model takes variable-length 5’UTR sequences as input to predict mean ribosome load (MRL). It combines a Saliency-Aware Token Clustering (SATC) module for selecting informative tokens with a Saliency-Guided Transformer (SGT) block for modeling regulatory interactions. The output representations are biologically interpretable, enabling recovery of known sequence elements.
Saliency-Aware Token Clustering Module
The adaptive tokenization process, implemented by the Saliency-Aware Token Clustering (SATC) module and illustrated in Figure.1, aggregates discrete input tokens into compact saliency-guided representations through two primary steps: (1) Saliency Gating, which identifies biologically informative tokens by assigning interpretable saliency scores and leverages these scores to guide subsequent clustering based on a density-distance potential; (2) Saliency-Preserving Centroid Synthesis, which aggregates the selected tokens within each cluster into unified representations via saliency-weighted averaging. This transformation effectively reduces token redundancy while preserving biologically informative features, resulting in more compact and interpretable embeddings that facilitate efficient modeling in subsequent layers.
Saliency Gating
Before any compression, we apply saliency gating to decide which tokens are worth keeping. Given a token vector $t_{i}\!\in\!\mathbb{R}^{d}$, and a learned weight vector $\mathbf{w}\!\in\!\mathbb{R}^{d}$, a single-head projection produces
$\alpha_{i}=\exp(\mathbf{w}^{\!\top}t_{i}).$ (1)
In practice, $\alpha_{i}$ acts as a interpretable saliency scores, indicating the relative importance of token $i$. These weights are later used as positional importance factors for biological motifs (e.g., Kozak consensus, uAUG upstream start codons), offering interpretability without additional heads.
With the interpretable saliency scores fixed, we measure token cohesion and separation via a density-distance potential. We first build a scale-normalised distance matrix
$D_{ij}=\frac{\|t_{i}-t_{j}\|_{2}}{\sigma\sqrt{d}},$ (2)
where $t_{i}\in\mathbb{R}^{d}$ denotes the $i$-th token embedding and $d$ is the feature dimension. The distance is normalized using the standard deviation of each embedding dimension, which reduces the dominance of high-variance features and improves the stability of distance computation. We then estimate the local density $\rho_{i}$ of each token using its $k$-nearest neighbors(Cover and Hart 1967):
$\rho_{i}=\exp\!\Bigl{(}-\tfrac{1}{k}\sum_{j\in\mathcal{N}_{k}(i)}D_{ij}^{2}\Bigr{)},$ (3)
where $\mathcal{N}_{k}(i)$ denotes the set of the $k$ nearest neighbors of token $i$. Then, for each token $i$, we explicitly compute a distance indicator $\delta_{i}$ representing the local density gap (Du, Ding, and Jia 2016) as:
$\delta_{i}=\begin{cases}\min_{j:\rho_{j}>\rho_{i}}D_{ij},&\text{if }\exists j:\rho_{j}>\rho_{i}\\ \max_{j}D_{ij},&\text{otherwise}\end{cases}$ (4)
where $\delta_{i}$ captures the minimal distance from token $i$ to any other token with a higher local density. For the token with the highest local density, $\delta_{i}$ is set to the maximum distance to any other token.
$\quad\gamma_{i}=\rho_{i}\,\delta_{i},$ (5)
where $\gamma_{i}$ represents the overall density–distance potential of token $i$ and guides the selection of representative centroids during clustering.
Model MPRA-U MPRA-H MPRA-V
[…]
RNA-FM 0.679 0.827 1.079 0.550 0.751 0.892 0.369 0.586 1.117
OptimusN 0.915 0.939 0.782 0.783 0.866 0.565 0.814 0.888 0.605
Framepooling 0.895 0.909 0.818 0.796 0.881 0.543 0.802 0.853 0.681
MTrans 0.941 0.954 0.447 0.812 0.902 0.526 0.845 0.909 0.552
UTR-LM 0.947 0.962 0.438 0.825 0.906 0.502 0.866 0.912 0.537
mRNA2vec 0.923 0.943 0.512 0.810 0.889 0.517 0.831 0.895 0.545
UTR-STCNet 0.956 0.967 0.388 0.849 0.914 0.471 0.907 0.937 0.429
Table 1: Performance Comparison of Single-Task Models Across Three 5’UTR Translation Datasets
Saliency-Preserving Centroid Synthesis
After computing the unified density-aware score $\gamma_{i}$, we select the top $\lceil\tau N\rceil$ tokens to serve as centroids, where $\tau\in(0,1)$ is a tunable hyperparameter denoting the proportion of selected tokens out of the total number $N$. The selection is based directly on the scores $\gamma_{i}$, which reflect each token’s contextual importance. This efficient, single-pass procedure identifies representative regions without requiring iterative refinement steps such as Expectation–Maximization (EM) or clustering post-processing.
Once clustering assignments are fixed, tokens within each cluster $c$ are aggregated via a saliency-weighted average:
$z_{c}=\sum_{i\in c}\left(\frac{\alpha_{i}}{\sum_{j\in c}\alpha_{j}}\right)t_{i},$ (6)
where $t_{i}$ is the contextual representation of token $i$ before clustering and $\alpha_{i}$ denotes the saliency scores obtained from the earlier Saliency Gating module (Eq. 1), explicitly capturing each token’s contextual significance.
By leveraging saliency as a soft weighting in the fusion step, the SATC module selectively emphasizes tokens with high functional or semantic relevance, while down-weighting less informative ones, resulting in more compact and discriminative representations.
Saliency Guided Transformer Block
The Saliency-Guided Transformer (SGT) block refines token representations by emphasizing biologically informative features and reducing redundancy. It models local and long-range dependencies through Saliency-Guided Aggregation, which compresses token features using saliency scores, preserving crucial contextual and structural information. Additionally, it integrates Motif-Level Contextualization via depth-wise convolutions that efficiently capture position-specific motifs critical for regulatory elements governing gene expression.
Saliency Guided Aggregation
The quadratic complexity of Transformer models poses a computational bottleneck for modeling long sequences, especially in 5’UTRs, where repetitive tokens often obscure essential regulatory information. To address this, we propose a Saliency-guided Aggregation (SGA) module, which reduces sequence length by aggregating Key and Value features according to their token-level saliency scores, thereby preserving structural and contextual information.
Specifically, given Query, Key, and Value matrices $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{B\times N\times C}$ and the saliency score matrix $\mathbf{P}\in\mathbb{R}^{B\times N\times 1}$ computed by the SATC module, we directly utilize these saliency scores as confidence weights to guide the subsequent local aggregation process.
For local aggregation, we partition the original sequences into segments $\Gamma_{\ell}$, each containing $r$ tokens, resulting in a reduced-length sequence with length $M=\lceil N/r\rceil$. Within each segment $\Gamma_{\ell}$, we denote the token-level Key and Value vectors at position $t$ as $\mathbf{k}_{t},\mathbf{v}_{t}\in\mathbb{R}^{C}$, and the scalar saliency score $p_{t}\in\mathbb{R}$, extracted from the saliency matrix $\mathbf{P}$. We then perform saliency-weighted aggregation as follows:
$\begin{split}&\widetilde{\mathbf{K}}_{\ell}=\frac{\sum_{t\in\Gamma_{\ell}}p_{t}\odot k_{t}}{\sum_{t\in\Gamma_{\ell}}p_{t}+\varepsilon},\\ &\widetilde{\mathbf{V}}_{\ell}=\frac{\sum_{t\in\Gamma_{\ell}}p_{t}\odot v_{t}}{\sum_{t\in\Gamma_{\ell}}p_{t}+\varepsilon},\\ \end{split}$ (7)
where $\varepsilon>0$ ensures numerical stability, and $\odot$ denotes element-wise multiplication. Stacking the aggregated vectors across all segments yields the reduced-length Key and Value matrices $\tilde{\mathbf{K}},\tilde{\mathbf{V}}\in\mathbb{R}^{B\times M\times C}$, preserving salient structural dependencies. Additionally, this method substantially reduces the spatial dimension of Key and Value, leading to decreased computational and memory overhead, and effectively models local continuity, thereby better preserving local token correlations.
In the subsequent attention layer, the clustered token representations serve as the Query $\mathbf{Q}$, while the compressed Key and Value representations ($\widetilde{\mathbf{K}},\widetilde{\mathbf{V}}$) are used as the context. To ensure that critical tokens maintain strong influence, we directly integrate token-level saliency scores into the attention logits by constructing an attention bias matrix $\mathbf{P}_{\mathrm{attn}}$. Then, the attention is computed as:
$\text{Attention}(\mathbf{Q},\widetilde{\mathbf{K}},\widetilde{\mathbf{V}})=\text{softmax}\left(\frac{\mathbf{Q}\widetilde{\mathbf{K}}^{\top}}{\sqrt{d_{k}}}+\mathbf{P}_{\mathrm{attn}}\right)\widetilde{\mathbf{V}}.$ (8)
This formulation enables high-saliency tokens to retain dominance in downstream attention, even after SGA, thereby achieving a balance between compression efficiency and information retention.
Model MPRA-H
[…]
w/o STCM & SGA 0.826 0.889 0.533
w/o SGA 0.837 0.905 0.486
w/o STCM 0.833 0.896 0.504
UTR-STCNet 0.849 0.914 0.471
Table 2: Effectiveness of different modules in UTR-STCNet.
Motif-Level Contextualization
Although the aforementioned SATC module compresses the sequence along its sequence dimension, the feed-forward layer in standard transformers remains strictly position-wise and therefore struggles to capture short-range compositional patterns such as 3 to 5 nucleotide motifs, which are critical for 5’UTR regulation. To address this, we redesign the feed-forward sub-layer by incorporating a depth-wise convolutional structure (Chollet 2017) that effectively captures both known and novel motif signals, as well as other functionally relevant sequence elements. Specifically, the revised module applies a 1D depth-wise convolution followed by a GELU activation(Hendrycks and Gimpel 2016), with a residual connection to ensure gradient stability. Depth-wise convolutions are particularly well-suited for modeling local dependencies in sequence data, as they apply channel-specific filters to capture fine-grained patterns with high parameter efficiency. The projection weights before and after the convolution are shared across tokens, preserving computational efficiency while enabling each channel to extract distinct local motif features.
This depth-wise convolutional mechanism introduces positional awareness at the motif level in a channel-efficient manner, complementing self-attention without requiring additional positional encodings. This mechanism further enables adaptation to varying input channel sizes following token aggregation, while preserving the spatial arrangement of tokens and capturing multi-scale features within the sequence. It integrates seamlessly with the channel-based token representations produced by the preceding SAG, enabling biologically meaningful context modeling into the SGT backbone. This enables the capture of functional differences across genes and complex regulatory contexts, enhances the modeling of dependencies between regulatory elements and sequence features, and ultimately improves the interpretability of functional signals in RNA sequences.
Figure 2: Prediction performance of UTR-STCNet on MRL. Evaluation was conducted on MPRA datasets with fixed-length sequences (left, n = 20,000) and variable-length sequences (right, n = 7,600). Performance was measured using the coefficient of determination ($\text{R}^{2}$) between observed and predicted MRL values on held-out test sets.
Experiments
Dataset
We evaluated the proposed UTR-STCNet model using three publicly available 5’UTR translation datasets: MPRA-U, MPRA-H, and MPRA-V. These datasets were all derived from massively parallel reporter assays (MPRAs)(Cuperus et al. 2017), providing abundant sequence information and corresponding expression levels to support model training and validation. We first ranked all sequences based on total read counts obtained from sequencing, and selected the top 20,000 most highly expressed sequences from the MPRA-U dataset as test samples. Similarly, the top 7,600 highly expressed human-derived sequences from the MPRA-H dataset were selected as the test set. For the MPRA-V dataset, due to the variability in sequence lengths, we adopted a stratified random sampling strategy: sequences were divided into length bins ranging from 25 to 100 nucleotides, and from each bin, 100 randomly generated sequences and 30 natural human sequences were selected, resulting in a total of 9,880 test sequences. The remaining unselected sequences from each dataset were used to construct the combined training and validation sets.
Baselines
We compare UTR-STCNet with the following seven state-of-the-art baseline models:
  • Optimus (Sample et al. 2019) employs convolutional networks with residual blocks to forecast mean ribosome load across fixed-length 5’ UTRs.
  • FramePool (Karollus, Avsec, and Gagneur 2021) is a length-agnostic CNN baseline that predicts 5’UTR mean ribosome load using frame pooling.
  • RNA-FM(Chen et al. 2022) is a foundation model that captures RNA structure, 3D proximity, and RBP interactions from unannotated sequences.
  • MTtrans (Zheng et al. 2023) shares a convolutional encoder across tasks while task‑specific GRU towers yield length‑agnostic, cross‑assay translation‑rate predictions.
  • UTR-LM (Chu et al. 2024) is a transformer-based language model that enables unified functional modeling and pretraining over 5’UTR sequences.
  • mRNA2vec(Zhang et al., 2025) is a general-purpose pretrained model based on data2vec jointly encoding 5’UTR and CDS with contextual and structural signals for mRNA representation.
Figure 3: Top enriched 3-mer motifs identified within high-saliency regions. High-saliency regions were first identified as stretches of consecutive nucleotides with elevated saliency scores. Within these regions, a 3-mer sliding window (step size = 3) was applied to extract all possible 3-mer motifs. A 3-mer was counted once per sequence if it appeared in a high-saliency region and exceeded a predefined score threshold. The top 15 3-mers with the highest frequency across all sequences are shown.
Performance Comparison
Single-Task Evaluation
Table 1 demonstrates that UTR-STCNet consistently achieves state-of-the-art performance across the three 5’UTR translation datasets: MPRA-U, MPRA-H, and MPRA-V. The model is comprehensively evaluated in terms of its generalization capability and predictive accuracy. On the MPRA-U dataset, UTR-STCNet achieves approximately 11.4% lower prediction error than UTR-LM, reflecting superior modeling accuracy.This result suggests that the model more effectively captures the complex nonlinear relationships between sequence features and expression levels, maintaining superior predictive performance even for sequences with high expression or strong regulatory signals. On the MPRA-H dataset, UTR-STCNet outperforms mRNA2vec by around 4.8%, highlighting its enhanced ability to model gene expression on real human-derived sequences. This performance gain demonstrates that UTR-STCNet is better equipped to learn intrinsic regulatory features and expression patterns in biologically authentic contexts, providing improved reliability and application potential in human transcriptome modeling. On the MPRA-V dataset, UTR-STCNet achieves a 3.1% improvement in Spearman correlation compared to MTrans. This dataset includes variable-length sequences, designed to rigorously evaluate the model’s robustness to input heterogeneity and its ability to capture long-range dependencies. The observed performance further validates UTR-STCNet’s architectural advantages in modeling complex contextual structures and its practical applicability to diverse 5’UTR length distributions found in real biological scenarios.
Figure 4: Distribution of saliency scores across TG-related groups. Density plots show the distribution of token-level saliency scores for different TG-related groups in panels a and b.
Multi-Task Evaluation
As shown in Figure 2, UTR-STCNet consistently achieves superior performance in multi-task settings on both fixed-length and variable-length synthetic MPRA datasets. On the fixed-length dataset (50 nt), UTR-STCNet effectively integrates multi-task information and achieves an $R^{2}$ score of 0.965, significantly outperforming the multi-task baseline model MTrans, demonstrating its stronger capability in modeling gene expression. On the more challenging variable-length dataset (25–100 nt), UTR-STCNet also outperforms all three baseline models, indicating its ability to dynamically adapt to input length variations and fully leverage contextual structures, thereby substantially enhancing prediction accuracy and modeling stability. Overall, UTR-STCNet not only delivers superior performance in multi-task scenarios, but also demonstrates greater robustness and generalization in handling variable-length inputs, which remain a fundamental challenge in real biological sequences, thereby highlighting its promising potential for practical applications.
Ablation Study
To evaluate the contribution of each core module in UTR-STCNet, we perform an ablation study on the MPRA-H dataset, as shown in Table 2. Specifically, we examine three ablation variants: removing both the Saliency-Aware Token Clustering Module STCM and the Saliency-Guided Aggregation Module SGA, removing only SGA while retaining STCM, and removing only STCM while keeping SGA. The results indicate that the combination of STCM and SGA yields the strongest performance, with an $R^{2}$ score of 0.849, Spearman correlation of 0.914, and RMSE of 0.471. In comparison, removing the SGA module alone reduces the $R^{2}$ to 0.837, showing that SGA plays a crucial role in preserving local continuity and guiding attention-weighted channel fusion. Similarly, removing STCM reduces the $R^{2}$ to 0.833, confirming that saliency-aware token clustering contributes to compressing redundant regions and enhancing structural coherence. When both modules are removed, the model performs worst, with $R^{2}$ dropping to 0.826 and RMSE increasing to 0.533. This highlights the limitations of directly applying vanilla Transformers to long UTR sequences without guided compression or context-aware aggregation. These findings validate the effectiveness of both modules. STCM improves input compactness through structural guidance, while SGA enhances downstream predictions by emphasizing important signals during aggregation. Together, they contribute synergistically to the model’s overall accuracy and generalization.
Biological Relevance of Model-Identified Regulatory Features
Figure 5: In silico mutagenesis reveals the functional impact of TG-related motifs on MRL predictions. a, Predicted MRL values after replacing the nucleotide preceding the ‘TG’ motif with adenine (A). Dark blue: mutated sequences; light blue: original experimental sequences. b, Predicted MRL values after replacing the ‘TG’ motif itself with non-TG variants. Dark blue: substituted sequences; light blue: original sequences.
Identification of High-Scoring Regions and Key Motifs
To investigate the regulatory features prioritized by UTR-STCNet, we designated high-scoring regions as contiguous stretches of three or more nucleotides with saliency scores exceeding 0.5, yielding a total of 22,393 candidate motifs. To systematically characterize local sequence patterns within these regions, we conducted a sliding window analysis using a 3-mer window. We computed the frequency of each 3-mer across all high-scoring regions and found that sequences containing the dinucleotide ‘TG’ were markedly overrepresented Figure 3, suggesting that ‘TG’-containing motifs may serve as important signals for MRL prediction. To validate this finding, we repeated the analysis using a 2-mer window. Consistent with the 3-mer results, ‘TG’ remained the most enriched 2-mer, reinforcing its role as a recurrent sequence feature learned by the model.
Impact of Adjacent Nucleotides to TG on MRL Prediction.
MRL Distributions of TG-Associated Nucleotide Combinations.
We analyzed the distribution of experimen tally measured MRL values for different TG-associated nu cleotide combinations. As shown in Figure 4a, sequences containing the ATG trinucleotide exhibited significantly lower MRL values compared to other combinations, suggesting that ATG may act as a translation-suppressive element (Andreev et al. 2022). In contrast, varying the nucleotide immediately following TG (A, C, G, or T) showed no substantial differences in MRL distributions Figure 4b, indicating that among trinucleotides containing TG, the presence of upstream ATG (uATG) is specifically associated with a negative regulatory effect (Tsotakos et al. 2015).
Figure 6: From the high-saliency regions, sequences centered on the dinucleotide ‘TG’ were extracted by taking 4 nucleotides upstream and 1 nucleotide downstream. The nucleotide frequencies flanking ‘TG’ were then computed to analyze base composition patterns surrounding this motif.
To validate the negative regulatory effect of ATG on MRL, we conducted an extreme test by replacing any non-A base immediately preceding each TG with A, while leaving other bases unchanged. This resulted in a significant overall decrease in MRL. This resulted in a significant overall decrease in MRL Figure 5a, further supporting the conclusion that ATG negatively impacts translation efficiency. Finally, to confirm whether the inhibitory effect is caused by the ATG triplet or the TG sequence itself, we randomly replaced all TGs with non-TG combinations Figure 5b. The prediction results indicated that while TG slightly reduced reaction efficiency, the primary inhibitory effect was attributed to the ATG sequence.
Feature Extraction Near Translation Initiation Sites.
To evaluate whether the model can capture regulatory patterns near translation initiation sites, we analyzed the nucleotide composition surrounding high-saliency TG motifs. For each TG site within high-scoring regions, we extracted four upstream and one downstream nucleotide and computed the positional nucleotide frequency distribution. The analysis revealed that the surrounding sequence context closely resembles the classical Kozak motif Figure 6 (Kozak 1986), indicating that the model effectively captures known regulatory features associated with translation efficiency.
Conclusion
In this work, we propose UTR-STCNet, a deep learning framework for modeling translational regulation from 5’UTR sequences. By combining a Saliency-Aware Token Clustering (SATC) module with a Saliency-Guided Transformer (SGT) block, UTR-STCNet captures long-range dependencies and key regulatory motifs with high interpretability. Extensive experiments demonstrate that the model consistently outperforms existing approaches across species and cell types, achieving state-of-the-art performance in mean ribosome load (MRL) prediction. In addition to accurate inference, UTR-STCNet reveals biologically meaningful patterns, recovering canonical elements such as uAUGs and Kozak motifs. These results highlight UTR-STCNet as a promising tool for advancing mRNA design and understanding translation control.